OpenStudy (anonymous):

what is the ratio of 5 miles to 35 miles in simplest form?

OpenStudy (larseighner):

5/35 = 1/7

OpenStudy (anonymous):

\[\Large \text{Welcome \to Openstudy}!!!!!\]

OpenStudy (anonymous):

1:7

sammixboo (sammixboo):

\(\large\sf\frac{5}{35}=\frac{1}{7}\)